OpenStudy (anonymous):

Ok, my algebra 2 teacher gave us this: f(x), f(-x), and -f(x) what do these mean?

hartnn (hartnn):

is it for even and odd functions ?

OpenStudy (anonymous):

no, he gave them to us with other letters like h(x) and g(x) and k(x) resembling compressions and stretches and functions. i just got them totally confused and dont know which ones are which

hartnn (hartnn):

oh...can you post a specific question so that i can help you in one certain direction.....this is more general

OpenStudy (anonymous):

Ok, one second

OpenStudy (anonymous):

For the functions f(x), write a function g(x) for the transformations: f(x)=-x-1; up 2 units, right 1 unit, compression by 1/4 <~~~ this is what he gave us on our worksheet

hartnn (hartnn):

ok, u have f(x)=-x-1 when you shift it to right by one unit you increase the x-cordinate by 1. so new x-coordinate X = x+1, did u get this?

OpenStudy (anonymous):

Ok, i understand that part of it

hartnn (hartnn):

now there is vertical shift of up by 2 so new Y = y+2 ok?

OpenStudy (anonymous):

ok, so left and right is increasing the X and up and down is increasing the Y?

hartnn (hartnn):

left is increasing the x right is decreasing the x up is increasing the y, down is decreasing the y

OpenStudy (anonymous):

Ohhhh, ok, ok, i got it!

hartnn (hartnn):

compression by 1/4 means if f(x) was 4 earlier, it becomes 1 now.

OpenStudy (anonymous):

Okkkk, is there some sort of formula for that? like the g(x) or k(x) or f(x) or h(x) with a + or - in front of them? i got really confused when my teacher was getting into that stuff

hartnn (hartnn):

read this: http://people.richland.edu/james/lecture/m116/functions/translations.html ask if any doubts.

OpenStudy (anonymous):

Oh, ok, that helps. thank you so much! :)

hartnn (hartnn):

welcome ^_^